OpenStudy (anonymous):

What is the solution to -2(3x - 9) + 5x = -10?

OpenStudy (anonymous):

Distribute the negative 2 and then combine like terms. Then solve for x algebraically. -2(3x - 9) + 5x = -10 -6x + 18 + 5x = -10 -x = -28 JUst divide both sides by -1 to get what x equals.